OpenStudy (anonymous):

Solve the triangle f a=4, b=3, and A=36(degrees) (find c, B, and C)

OpenStudy (yttrium):

Use cosine law: a^2 = b^2 + c^2 - 2bc cosA Isolate c, and solve for c. After solving for C, you can easily solve for the remaining angles by using cosine or sine law. Depends on your choice.

OpenStudy (anonymous):

I don't understand how to work it out. Can you walk me through the steps?

OpenStudy (yttrium):

Where are you stucked?

OpenStudy (anonymous):

\[c ^{2}=4^{2}+3^{2}-2*4*3\cos36 ... that is the formula but how do I evaluate \it? Pemdas?\]

OpenStudy (anonymous):

like, I think exponents, then the basic math. Then I have c^2=1cos36... thats not right

OpenStudy (yttrium):

Well it will become: c^2 - 6c cos 36 = 4^2 - 3^2 c^2 - 6c cos 36 = 7 c^2 - 6c cos 36 - 7 = 0 Apply quadratic formula to solve for C.